OpenStudy (anonymous):

Richard rides a bike on a road sloping downward and inclined at an angle of 30 degrees with the horizontal. If he moves at a constant speed of 10meters/second, calculate his vertical downward velocity

OpenStudy (ivancsc1996):

Ok so what you need to do is break the velocity "v" into its components "vy" and "vx".They are asking you specifically for the component "vy" since it is vertical or in other owrds in the "y" direction. If you have seen a little bit of trigonometry you will se that the sine of 30° is equal to the component "vy" divided by the velocitty "v". If you solve the for "vy" you get the equation you need. The you just need to plug 10meters/second in "v" and you get the answer.\[\sin(30°)=\frac{ v _{y} }{ v }\rightarrow v _{y}=vsin(30°)=10\frac{ m }{ s }\sin(30°)=5\frac{ m }{ s }\]